A manufacturer maintains a supply of a key part used in the production of their product by purchasing the part from two different suppliers A and B. Because this is a key component and they want to maintain a good relationship with both suppliers, the warehouse inventory is monitored regularly. Currently there are 854 unit parts in the warehouse of which 472 more are from supplier A than there are from supplier B. How many in the warehouse are from supplier B?
Answer by jainenderkapoor(61)


Let there are `x` quantities from supplier B,
then there are `x` + 472 quantities from supplier A

we get x + x + 472 = 854
2x = 854 - 472
2 x = 382
x = 382/2
x = 191
So 191 units are from supplier `B`.


I hope that the solution is clear to you.
If you have any doubt you are welcome to contact me.
I am also available for online tutoring in math at very reasonable rates.
My mail id is jainenderkapoor@gmail.com
Jainender Kapoor